OBSERVATION OF THE 5 ^{1}Σ_{u}^{+} and 5 ^{1}Π_{u} STATES OF Rb_{2} BY POLARISATION LABELLING SPEC-TROSCOPY ^{a}

JACEK SZCZEPKOWSKI, WŁODZIMIERZ JASTRZȨBSKI, Institute of Physics, Polish Academy of Sciences, Al. Lotnikóv 32/46, 02-668 Warsaw, Poland ; PAWEŁ KOWALCZYK, ANNA GRO-CHOLA, Institute of Experimental Physics, Department of Physics, University of Warsaw, ul. Hoża 69, 00-681 Warsaw, Poland ; ABDUL-RAHMAN ALLOUCHE, PATRICK CROZET and \underline{AMANDA J. ROSS}, Institut Lumière Matière, Université Lyon 1 & CNRS UMR5306, Université de Lyon, France.

Two-colour polarization labeling experiments have been used to explore the excitation spectrum of the rubidium dimer in the region 25500 - 27000 cm^{−1}, aiming to establish the symmetry of excited states observed by Han and Heaven^{b} in earlier work. The spectra we recorded in fact probe quite different electronic states; we identify them from ab initio calculations as the 5 ^{1}Σ_{u}^{+} and 5 ^{1}Π_{u} states, dissociating to Rb 5s + Rb 5d atoms. The home-made pulsed laser used to label levels in the electronic ground state, with nominal linewidth of 0.3 cm^{−1}, is not sufficiently selective for this heavy-alkali species. The spectra are a tangled confusion of short and rather irregular progressions of P,R doublets and/or Q lines. We have had to use cross-correlation techniques to identify transitions seen in different spectra that share a common ground state level, to establish assignments. Analysis confirms unusual vibrational patterns in the lowest vibrational levels, from avoided crossings with ion pair states.

aSupported by the PAN-CNRS exchange programme, 2011-12

bJ. Han, M. Heaven, J. Mol. Spectrosc. \underline{268} 37-41 (2011)